# FiRe: Fixed-points of Restoration Priors for Solving Inverse Problems

### v1

@@ 6 Conclusion
We have introduced Fixed-points of Restoration (FiRe) priors, extending the implicit prior paradigm beyond denoising to general restoration models. Our approach leverages the observation that natural images are fixed points of restoration models composed with their training degradation operators. This yields both an explicit prior formulation and a versatile algorithm requiring no restrictive assumptions on the degradation prior. Experiments across various restoration tasks demonstrate two key capabilities: the combination of multiple restoration models in an ensemble-like fashion, and the incorporation of measurement-aware priors. These features enable FiRe to address challenging scenarios where traditional PnP methods struggle, such as noisy inpainting, all within a unified framework.
Acknowledgments
This work was supported by the BrAIN grant (ANR-20-CHIA-0016) and was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014344 made by GENCI.
@@ Appendix A Convergence of stochastic algorithm
First, we have that in our setting, $f$ is convex. Moreover, we have that $d_{C}^{2}$ is bounded from below, and from Section 3.1, we have that $\nabla d_{C}^{2}$ is Lipschitz. The result then follows from [18, Theorem Corollary 3.6]. ∎

### v2

@@ 6 Conclusion
We have introduced Fixed-points of Restoration (FiRe) priors, extending the implicit prior paradigm beyond denoising to general restoration models. Our approach leverages the observation that natural images are fixed points of restoration models composed with their training degradation operators. This yields both an explicit prior formulation and a versatile algorithm requiring no restrictive assumptions on the degradation prior. Experiments across various restoration tasks demonstrate two key capabilities: the combination of multiple restoration models in an ensemble-like fashion, and the incorporation of measurement-aware priors. These features enable FiRe to address challenging scenarios where traditional PnP methods struggle, such as noisy inpainting, all within a unified framework. A challenge remains with the tuning of hyperparameters, which can significantly impact performance across different degradation settings. Addressing this limitation through adaptive or learned parameter selection would be a promising direction for future research.
Acknowledgements
This work was supported by the BrAIN grant (ANR-20-CHIA-0016) and was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014344 made by GENCI. Ulugbek Kamilov was supported by the NSF CAREER award under CCF-2043134.
Supplementary Material
@@ Appendix A Convergence of stochastic algorithm
First, we have that in our setting, $f$ is convex. Moreover, we have that $d_{C}^{2}$ is bounded from below, and from Section 3.1, we have that $\nabla d_{C}^{2}$ is Lipschitz. The result then follows from [19, Theorem Corollary 3.6]. ∎
@@ C.2 Inpainting models
Appendix D Additional visual results
Appendix E Influence of the degradation
The choice of degradation operator significantly impacts reconstruction quality. In Fig. 10, we give reconstruction metrics for varying degradation strengths using either a Gaussian Restormer prior and a DRUNet denoising prior. This shows that a minimum degradation is required to stabilize the reconstruction, while increasing beyond a certain threshold leads to excessive smoothing.
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Influence of the degradation operator on the FiRe restoration algorithm on the Set3C algorithm for different problems. Left plot: FiRe with Restormer Gaussian prior. We run the algorithm with different kernel widths. Reconstructions (details) corresponding to points (b), (c), (d) on the plot are shown on the left. Right plot: same, but with a DRUNet denoiser restoration prior; (f), (g) and (h) panels show the associated reconstructions. (a) and (e) show the degraded measurements $y$.